# Cuff-less Arterial Blood Pressure Waveform Synthesis from Single-site PPG using Transformer & Frequency-domain Learning

### v1

@@ Abstract
We propose two novel purpose-built deep learning (DL) models for synthesis of the arterial blood pressure (ABP) waveform in a cuff-less manner, using a single-site photoplethysmography (PPG) signal. We utilize the public UCI dataset on cuff-less blood pressure (CLBP) estimation to train and evaluate our DL models. Firstly, we implement a transformer model that incorporates positional encoding, multi-head attention, layer normalization, and dropout techniques, and synthesizes the ABP waveform with a mean absolute error (MAE) of 14. Secondly, we implement a frequency-domain (FD) learning approach where we first obtain the discrete cosine transform (DCT) coefficients of the PPG and ABP signals corresponding to two cardiac cycles, and then learn a linear/non-linear (L/NL) regression between them. We learn that the FD L/NL regression model outperforms the transformer model by achieving an MAE of 11.87 and 8.01, for diastolic blood pressure (DBP) and systolic blood pressure (SBP), respectively. Our FD L/NL regression model also fulfills the AAMI criterion of utilizing data from more than 85 subjects, and achieves grade B by the BHS criterion.
@@ I Introduction
The existing clinical-grade methods for blood pressure estimation (including ambulatory devices) could be classified into two categories: i) invasive methods, ii) non-invasive but cuff-based methods. Invasive methods are the gold standard methods which measure the centralized aortic blood pressure by inserting catheters in the aorta (main artery near the heart) [6]. Need not to say that this method is painful, prone to infections, and required trained professionals in a clinical setting. Next, non-invasive but cuff-based methods work by inflating the cuff around the arm and rely upon manual auscultation and Sphygmomanometers for peripheral blood pressure estimation. These methods, though highly accurate, suffer from a number of limitations, e.g., inconvenience of regular monitoring, requirement of trained personnel, Sphygmomanometers use toxic mercury, etc [7, 8, 9]. Further, the need for frequent cuff inflation and deflation makes continuous BP measurement impractical, especially for individuals with weaker arms, or the elderly [10]. Oscillometric devices, though do not require manual auscultation, have relatively lower accuracy as they rely on mean arterial blood pressure (MABP) for estimating both SBP and DBP.
The limitations of the existing invasive methods and cuff-based non-invasive methods have thus prompted the researchers to design novel cuff-less blood pressure (CLBP) estimation methods which are meant to be user-friendly, reliable, and support continuous and at-home BP measurement [7, 11, 12, 13]. Additionally, advancements in wearable technology (e.g., smart watches, smart bands, fitness trackers, internet of medical things, flexible electronic tattoos, etc.) and assistive gadgets (e.g., smartphones, airpods, etc.) have further accelerated the development of the CLBP estimation methods [14], [15]. CLBP estimation methods either rely upon bio-signals such as photoplethysmography (PPG), electrocardiography (ECG), phonocardiography (PCG) acquired non-invasively from one or more body sites [12], or innovative sensor technology, e.g., tonometry [16].
@@
  • We implement a transformer model that incorporates positional encoding, multi-head attention, layer normalization, and dropout techniques, and synthesizes the ABP waveform with a mean absolute error (MAE) of 14.
  • We implement a frequency-domain (FD) learning approach where we first obtain the discrete cosine transform (DCT) coefficients of the PPG and ABP signals corresponding to two cardiac cycles, and then learn a linear/non-linear (L/NL) regression between them.
The transformer model (the FD L/NL regression model) achieves an MAE of 19 and 9 (11.87 and 8.01), for diastolic blood pressure (DBP) and systolic blood pressure (SBP), respectively. Our FD L/NL model achieves grade B by the British Hypertension Society (BHS) criterion.
@@ II Background & Related Work
II-A Background
Pulse wave analysis methods: During the last decade, researchers have made extensive investigation of the physiological waveforms acquired in a non-invasive manner (e.g., ECG, PPG, PCG, etc.), under the umbrella term of pulse wave analysis (PWA) [12]. The aim of PWA methods is to develop a number of indices to indirectly measure the blood pressure in a cuff-less manner using wearables (e.g., smart watches) and assistive gadgets (e.g., smartphones) [14]. Some prominent indices (markers for blood pressure) include the following: pulse transit time (PTT), pulse arrival time (PAT), and pulse wave velocity (PWV). Before we delve into the discussion of the related work, it is imperative to formally define the three indices (BP surrogates) as follows:
@@ II-A Background
Despite the good correlation of PTT, PAT and PWV with the blood pressure, these indices have a number of limitations as well, e.g., need for simultaneous and synchronous measurement of two waveforms from two different sites, calibration issues, lack of consensus on translation of PTT/PAT/PWV to BP (i.e., there exist a number of equations to map PTT/PAT/PWV to BP, each with many unknowns that are subject-specific), etc. [21], [12], [19]. This highlights the need for the design of novel single-site methods for cuff-less blood pressure (CLBP) estimation.
@@ II-B Related work
Since the literature on CLBP estimation is quite vast, we could only summarize selected related works due to space constraints (see [12], [19] and the references therein, which provide a more detailed treatment of the subject). The CLBP estimation methods are mostly PWA methods which are either two-sites/two-waveforms methods or single-site methods. More recently, researchers have extensively used various deep learning (DL) models for ABP waveform synthesis and CLBP estimation.
@@ III-B Data Preprocessing
In order to prepare the data for our DL models, we implemented the following data preprocessing steps on the PPG and ABP signals. 1) We began by removing out-of-band noise from the data using a Butterworth low-pass filter with a cut-off frequency of 10 Hz. 2) We then proceeded to remove the residual artifacts manually (by discarding the anomalous samples). 3) Next, we removed the baseline from the data using the Wavelet filtering. 4) We then normalized the data using z-score normalization. 5) Further, the process of ABP waveform synthesis from PPG waveform required beat-level synchronization between the PPG and ABP signals. To this end, we utilized time shifting approach to align the systolic peaks of the PPG signal with the systolic peaks of the ABP signal. 6) We then did segmentation (with a stride of 0) to construct a number of segments from each PPG and ABP recording, in order to increase the size of the dataset. Note that each segment consists of two cardiac cycles.
@@ IV Methodology
Keeping in mind that the problem at hand—ABP waveform synthesis from PPG waveform—is a sequence-to-sequence translation (i.e., regression) problem, we have implemented two novel deep learning methods using Tensorflow and Keras frameworks in Python: 1) a transformer model, 2) a linear/non-linear frequency-domain regression model. Once the ABP waveform is synthesized, we sample its maximum and minimum values which yield the systolic and diastolic blood pressure, respectively. Below, we discuss the two methods in detail, one by one.
@@ IV-A Transformer model for ABP waveform synthesis
Our transformer model implementation consists of the following discrete components: positional encoding, multihead self-attention, feedforward network, layer normalization, dropout, and a dense output layer. Fig. 1 shows the architecture of the transformer implemented in this work in detail.
Figure 1: Architecture of Transformer model for ABP waveform synthesis.
At this point, it is imperative to revisit the individual building blocks of the transformer model to see how each component helps solve the problem at hand. 1) First, the positional encoding (often implemented by means of sine and cosine functions) keeps the transformer informed about the order and position of data points (i.e., the notion of time) in each of the two (i.e., PPG and ABP) sequences. 2) Next, the multi-head self-attention mechanism, as the name implies, focuses on capturing relationship between every pair of data points (by means of attention scores) in both the PPG and ABP sequences, using multiple heads that attend to different parts of the PPG and ABP sequences in order to capture the short term and the long term dependencies within the PPG and ABP signals. 3) Next, the layer Normalization reduces the internal covariate shift, prevents the issue of vanishing or exploding gradients, helps prevent overfitting, all by normalizing the activations. 4) The dropout regularization layer randomly sets a fraction of the values to zero during training, which helps to prevent overfitting. 5) Element-wise addition between the original input and the output of the dropout layer is performed in the next block. This residual connection allows information to bypass the self-attention and normalization layers, helping the model retain valuable information from the original input. 6) Next, the feedforward network applies a set of linear transformations and non-linear activations to the input data, allowing the model to learn complex patterns and relationships. 7) Finally, the dense layer produces the output of the model.
IV-B Linear/non-linear frequency-domain regression model for ABP waveform synthesis
@@
  • Frequency-domain representation of pre-processed data: Firstly, we obtain the frequency-domain representation of the pre-processed data, for both signals of interest, i.e., PPG and ABP signals. The motivation for this comes from the fact that both PPG and ABP signals are quasi-periodic due to the rhythmic activity of the heart. Specifically, we consider one PPG-ABP segment-pair corresponding to a single cardiac cycle of duration $Q^{\prime}$ samples, at a time. We then compute the discrete cosine transform (DCT) (type-II) of each of the PPG and ABP segments in the single PPG-ABP segment-pair. Further, it is well-known that the PPG and ABP signals are sparse in frequency domain (with only a few low-frequency components with significant energy). This allows us to have a compact representation of each of the two segments in the given PPG-ABP segment-pair by retaining first $Q_{X}$ ($Q_{Y}$) significant DCT coefficients of the PPG (ABP) segment only. This in turn helps in reducing the computationally complexity of our linear/non-linear (L/NL) regression model by reducing the dimension of the feature vector it takes as input. Note that we zero-pad both DCT vectors of size $Q_{X}$ and $Q_{Y}$ to resize them to $Q$ in order to do the L/NL regression in the next step (where $Q$ represents the duration of a cardiac cycle, in samples).
[…]
Next, the testing phase. We compute DCT (type-II) of a previously unseen size-$Q^{\prime}$ PPG segment, resize it to $Q_{X}$ and then $Q$. We then feed the size-$Q$ DCT vector (of PPG) to the L/NL frequency-domain regression model which outputs a size-$Q$ DCT vector (of ABP). Then, ABP waveform is synthesized by taking the inverse DCT (I-DCT) of the output of the L/NL regression model. Fig. 2 provides a pictorial overview of the proposed FD L/NL regression model.

### v2

@@ Abstract
We develop and evaluate two novel purpose-built deep learning (DL) models for synthesis of the arterial blood pressure (ABP) waveform in a cuff-less manner, using a single-site photoplethysmography (PPG) signal. We train and evaluate our DL models on the data of 209 subjects from the public UCI dataset on cuff-less blood pressure (CLBP) estimation. Our transformer model consists of an encoder-decoder pair that incorporates positional encoding, multi-head attention, layer normalization, and dropout techniques for ABP waveform synthesis. Secondly, under our frequency-domain (FD) learning approach, we first obtain the discrete cosine transform (DCT) coefficients of the PPG and ABP signals, and then learn a linear/non-linear (L/NL) regression between them. The transformer model (FD L/NL model) synthesizes the ABP waveform with a mean absolute error (MAE) of 3.01 (4.23). Further, the synthesis of ABP waveform also allows us to estimate the systolic blood pressure (SBP) and diastolic blood pressure (DBP) values. To this end, the transformer model reports an MAE of 3.77 mmHg and 2.69 mmHg, for SBP and DBP, respectively. On the other hand, the FD L/NL method reports an MAE of 4.37 mmHg and 3.91 mmHg, for SBP and DBP, respectively. Both methods fulfill the AAMI criterion. As for the BHS criterion, our transformer model (FD L/NL regression model) achieves grade A (grade B).
@@ I Introduction
The existing clinical-grade methods for blood pressure estimation (including ambulatory devices) could be classified into two categories: i) invasive methods, ii) non-invasive but cuff-based methods. Invasive methods are the gold standard methods which measure the centralized aortic blood pressure by inserting catheters in the aorta (main artery near the heart) [6]. Need not to say that this method is painful, prone to infections, and requires trained professionals in a clinical setting. Next, non-invasive but cuff-based methods work by inflating the cuff around the arm and rely upon manual auscultation and Sphygmomanometers for peripheral blood pressure estimation. These methods, though highly accurate, suffer from a number of limitations, e.g., inconvenience of regular monitoring, requirement of trained personnel, use of toxic mercury by Sphygmomanometers, etc. [7, 8, 9]. Further, the need for frequent cuff inflation and deflation makes continuous BP measurement impractical, especially for individuals with weaker arms, or the elderly [10]. Oscillometric devices do not require manual auscultation, but they have relatively lower accuracy as they rely on mean arterial blood pressure (MABP) for estimating both SBP and DBP.
The limitations of the existing invasive methods and cuff-based methods have thus prompted the researchers to design novel cuff-less blood pressure (CLBP) estimation methods which are meant to be user-friendly, reliable, and support continuous and in-situ BP measurement [7, 11, 12, 13]. Additionally, advancements in wearable technology (e.g., smart watches, smart bands, fitness trackers, internet of medical things, flexible electronic tattoos, etc.) and assistive gadgets (e.g., smartphones, airpods, etc.) have further accelerated the development of the CLBP estimation methods [14], [15]. CLBP estimation methods either rely upon bio-signals such as photoplethysmography (PPG), electrocardiography (ECG), phonocardiography (PCG) acquired non-invasively from one or more body sites [12], or innovative sensor technology, e.g., tonometry [16].
@@
  • We implement a transformer based method that incorporates positional encoding, multi-head attention, layer normalization, and dropout techniques, and synthesizes the ABP waveform with a mean absolute error (MAE) of 3.01. Further, the MAE of SBP and DBP are 3.77 and 2.69, respectively. This method satisfies AAMI criterion, and achieves grade A, according to BHS criterion.
  • We implement a frequency-domain (FD) learning method where we first obtain the discrete cosine transform (DCT) coefficients of the PPG and ABP signals, and then learn a linear/non-linear (L/NL) regression between them. This method synthesizes the ABP waveform with an MAE of 4.23. Further, the MAE of SBP and DBP are 4.37 and 3.91, respectively. This method satisfies AAMI criterion, and achieves grade B, according to BHS criterion.
We note that this work utilizes the UCI dataset (that contains data of patients in ICU), thus, the proposed methods do CLBP estimation under the condition when the subject is at rest.
@@ II Background & Related Work
Before we summarize the selected related work, it is imperative to briefly introduce the so-called pulse wave analysis methods, which act as scaffolding for the design of most of the CLBP estimation methods.
II-A Background: Pulse wave analysis methods
During the last decade, researchers have made extensive investigation of the physiological waveforms acquired in a non-invasive manner (e.g., ECG, PPG, PCG, etc.), under the umbrella term of pulse wave analysis (PWA) [12]. The aim of PWA methods is to develop a number of indices to indirectly measure the blood pressure in a cuff-less manner using wearables (e.g., smart watches) and assistive gadgets (e.g., smartphones) [14]. Some prominent indices (markers for blood pressure) include the following: pulse transit time (PTT), pulse arrival time (PAT), and pulse wave velocity (PWV). Before we delve into the discussion of the related work, it is imperative to formally define the three indices (BP surrogates) as follows:
@@ II-A Background: Pulse wave analysis methods
Despite the good correlation of PTT, PAT and PWV with the blood pressure, these indices have a number of limitations as well, e.g., need for simultaneous and synchronous measurement of two waveforms from two different sites. Furthermore, there is a lack of consensus on translation of PTT/PAT/PWV values to SBP and DBP values. For example, one way to measure the SBP and DBP values through PTT and PWV is as follows:
$\text{SBP}=\alpha_{\text{s}}+\beta_{\text{s}}\cdot\text{PTT}+\gamma_{\text{s}}\cdot\text{PWV}$ (1)
$\text{DBP}=\alpha_{\text{d}}+\beta_{\text{d}}\cdot\text{PTT}+\gamma_{\text{d}}\cdot\text{PWV}$ (2)
where the specific coefficients $\alpha_{\text{s}},\beta_{\text{s}},\gamma_{\text{s}},\alpha_{\text{d}},\beta_{\text{d}},\gamma_{\text{d}}$ need to be determined through empirical calibration with actual blood pressure measurements. Thus, such translation methods have calibration issues, as one needs to re-celibrate on new datasets. In fact, there exist a number of equations to map PTT/PAT/PWV to SBP and DBP values, each with many unknowns that are subject-specific [24], [12], [22].
This highlights the need for the design of novel single-site, calibration-free, non-parametric methods for cuff-less blood pressure estimation.
@@ II-B Related work
Since the literature on CLBP estimation is quite vast, we could only summarize selected related works due to space constraints (see [12], [22] and the references therein, which provide a more detailed treatment of the subject). The CLBP estimation methods are mostly PWA methods which are either two-sites/two-waveforms methods or single-site methods. More recently, researchers have extensively used various deep learning models for ABP waveform synthesis and CLBP estimation.
@@ III-B Data Preprocessing
In order to prepare the data for our DL models, we implemented the following data preprocessing steps on the PPG and ABP signals. 1) We began by removing out-of-band noise from the data using a Butterworth low-pass filter with a cut-off frequency of 10 Hz. 2) We then proceeded to remove the residual artifacts manually (by discarding the anomalous samples). 3) Next, we removed the baseline from the data by means of five-level Wavelet decomposition and reconstruction using db4 mother wavelet. 4) Further, the process of ABP waveform synthesis from PPG waveform required beat-level synchronization between the PPG and ABP signals. To this end, we utilized time shifting (cross-correlation) approach to align the systolic peaks of the PPG signal $x[n]$ with the systolic peaks of the ABP signal $y[n]$:
$R_{xy}(m)=\sum_{n=0}^{N-1}x[n]y[n-m]$ (3)
where $m$ is the lag variable, and ranges from $-N+1$ to $N-1$.
5) Inline with previous works [38, 39], we did segmentation (with a stride of 0) to construct a number of segments from each PPG and ABP recording, in order to increase the size of the dataset. Note that each segment is of duration 2 seconds (i.e., it consists of 2-3 cardiac cycles). 6) Finally, we normalized the segmented data using z-score normalization: $z=\frac{q_{s}-\mu}{\sigma}$. where $z$ is the normalized data, $x$ is the original data, $\mu$ is the mean across the data segment, and $\sigma$ is the standard deviation across the segment.
@@ IV Methodology
Keeping in mind that the problem at hand—ABP waveform synthesis from PPG waveform—is a sequence-to-sequence translation (i.e., regression) problem, we have implemented two novel deep learning methods using Tensorflow and Keras frameworks in Python: 1) a purpose-built transformer model, 2) a frequency-domain linear/non-linear regression model. Once the ABP waveform is synthesized, we sample its maximum and minimum values which yield the systolic and diastolic blood pressure, respectively.
@@ IV-A Transformer model for ABP waveform synthesis
Since we aim to solve a sequence-to-sequence translation problem, our purpose-built transformer model consists of an encoder-decoder pair. During the implementation phase, we tried several configurations of the transformer model, and saved the best performing configuration. Table I summarizes the architecture of our purpose-built transformer. One can see that both the encoder and decoder utilize a dense layer, followed by positional encoding layers, followed by three transformer blocks. The outputs of the decoder and encoder blocks are fused by means of a multi-head cross attention mechanism. This is followed by a couple of dense layers at the end which produce the final output, i.e., ABP signal.
Note that each transformer block consists of the following layers: multi-head self-attention mechanism and feed-forward neural networks, with residual connections and layer normalization applied after each sub-layer. Further, it is imperative to explain the working of the following building blocks of our transformer model:
1) Positional encoding: It keeps the transformer informed about the order and position of data points (i.e., the notion of time) in each of the two (i.e., PPG and ABP) sequences. For a position $pos$ and dimension $i$:
$PE_{(pos,2i)}=\sin\left(\frac{pos}{10000^{\frac{2i}{d_{model}}}}\right)\quad(\text{for even indices})$ (4)
$PE_{(pos,2i+1)}=\cos\left(\frac{pos}{10000^{\frac{2i}{d_{model}}}}\right)\quad(\text{for odd indices})$ (5)
where $d_{model}$ is the dimension of the model (i.e., the size of the embedding vectors).
2) Multi-head self-attention: As the name implies, it focuses on capturing relationship between every pair of data points (by means of attention scores) in both the PPG and ABP sequences, using multiple heads that attend to different parts of the PPG and ABP sequences in order to capture the short term and the long term dependencies within the PPG and ABP signals. Multi-head self-attention mechanism consists of the following key steps. Let $X$ represent input embeddings of shape $(n,d_{\text{model}})$ with $n$ being the sequence length. First, the query, key, and value matrices $Q,K,V$ are obtained by linear projections of the input embeddings as follows: $Q=XW_{Q}$, $K=XW_{K}$, $V=XW_{V}$ where $W_{Q}\in\mathbb{R}^{d_{\text{model}}\times d_{k}},\quad W_{K}\in\mathbb{R}^{d_{\text{model}}\times d_{k}},\quad W_{V}\in\mathbb{R}^{d_{\text{model}}\times d_{v}}$ are the three learned weight matrices, and $d_{k},d_{v}$ represent the dimension of the key, value vectors, respectively. This allows us to compute the scaled dot-product attention as follows:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (6)
where softmax is an activation function defined as follows: $\text{softmax}(\mathbf{z})_{i}=\frac{e^{z_{i}}}{\sum_{j=1}^{N}e^{z_{j}}}$ where $\mathbf{z}$ is the input vector, $z_{i}$ is the $i$-th element of the input vector $\mathbf{z}$, $N$ is the number of elements in the input vector $\mathbf{z}$. Finally, multi-head attention is computed as follows:
$\text{MultiHead}(Q,K,V)=\text{Concat}(\text{head}_{1},\ldots,\text{head}_{h})W^{O}$ (7)
where $h$ is number of heads, and $W^{O}$ is the output projection matrix. Further,
$\text{head}_{i}=\text{Attention}(Q_{i}=XW_{Q_{i}},K_{i}=XW_{K_{i}},V_{i}=XW_{V_{i}})$ (8)
where $W_{Q_{i}},W_{K_{i}},W_{V_{i}}$ are the query, key, value projection matrices for the i-th head.
Layer type Output Shape No. of Param Connected to Input (e) (None, 250, 1) 0 - Input (d) (None, 250, 1) 0 - Dense (e) (None, 250, 64) 128 Input (e) Dense (d) (None, 250, 64) 128 Input (d) Pos. Enc. (e) (None, 250, 64) 0 Dense (e) Pos. Enc. (d) (None, 250, 64) 0 Dense (d) Tx Block 1 (e) (None, 250, 64) 74944 Pos. Enc. (e) Tx Block 1 (d) (None, 250, 64) 74944 Pos. Enc. (d) Tx Block 2 (e) (None, 250, 64) 74944 Tx Block 1 (e) Tx Block 2 (d) (None, 250, 64) 74944 Tx Block 1 (d) Tx Block 3 (e) (None, 250, 64) 74944 Tx Block 2 (e) Tx Block 3 (d) (None, 250, 64) 74944 Tx Block 2 (d) MultiHead Attention (None, 250, 64) 66368 Tx Block 3 (e), Tx Block 3 (d) Dense 1 (None, 250, 64) 4160 MultiHead Attention Dense 2 (None, 250, 1) 65 Dense 1
TABLE I: Architecture of our purpose-built transformer model for ABP waveform synthesis. (e) implies an encoder layer, while (d) implies a decoder layer. Finally, Tx is the short for transformer.
IV-B Frequency-domain linear/non-linear regression model for ABP waveform synthesis
@@
  • Frequency-domain representation of pre-processed data: Firstly, we obtain the frequency-domain representation of the pre-processed data, for both signals of interest, i.e., PPG and ABP signals. The motivation for this comes from the fact that both PPG and ABP signals are quasi-periodic due to the rhythmic activity of the heart. Specifically, we consider one PPG-ABP segment-pair corresponding to 2-3 cardiac cycles of duration $Q^{\prime}$ samples, at a time. We then compute the discrete cosine transform (DCT) (type-II) of the PPG segment $x_{n}$ of length $Q^{\prime}$ in the single PPG-ABP segment-pair:
    $X_{k}=\sum_{n=0}^{Q^{\prime}-1}x_{n}\cos\left(\frac{\pi}{Q^{\prime}}\left(n+\frac{1}{2}\right)k\right)\quad k=0,\ldots,Q^{\prime}-1$ (9)
    Similarly, we compute the DCT of the corresponding ABP segment in the single PPG-ABP segment-pair. We observe that the PPG and ABP signals are sparse in frequency domain (with only a few low-frequency components with significant energy). This allows us to have a compact representation of each of the two segments in the given PPG-ABP segment-pair by retaining first $Q_{X}$ ($Q_{Y}$) significant DCT coefficients of the PPG (ABP) segment only. This in turn helps in reducing the computationally complexity of our linear/non-linear (L/NL) regression model by reducing the dimension of the feature vector it takes as input. Note that we zero-pad both DCT vectors of size $Q_{X}$ and $Q_{Y}$ to resize them to $Q$ in order to do the L/NL regression in the next step (where $Q$ represents the duration of a segment, in samples).
  • Linear/non-linear regression: Secondly, we learn the linear/non-linear regression between the size-$Q$ DCT coefficients vector $\mathbf{x}$ (that represents the most important features of the PPG signal) and the size-$Q$ DCT coefficients vector $\mathbf{y}$ (that represents the most important features of the ABP signal). For linear regression, we have implemented ridge regression [^2] Ridge regression, also known as Tikhonov regularization, is basically linear regression but with an additional term for $L2$ regularization that helps prevent overfitting and stabilize the regression coefficients.. The ridge regression solution is given by:
    $\mathbf{w}=(\mathbf{X}^{T}\mathbf{X}+\lambda\mathbf{I})^{-1}\mathbf{X}^{T}\mathbf{y}$ (10)
    where $\mathbf{X}$ is a $Q\times Q$ matrix of DCT vectors $\mathbf{x}$ of PPG segments, $\mathbf{y}$ is a $Q\times 1$ vector of DCT coefficients of ABP segment, $\lambda$ is a scalar regularization parameter, $\mathbf{I}$ is the identity matrix of size $Q\times Q$, $\mathbf{w}$ is the vector of regression coefficients.
Next, the testing phase. We compute DCT (type-II) of a previously unseen size-$Q^{\prime}$ PPG segment, resize it to $Q_{X}$ and then $Q$. We then feed the size-$Q$ DCT vector (of PPG) to the L/NL frequency-domain regression model which outputs a size-$Q$ DCT vector (of ABP). Then, the ABP waveform $x_{n}$ is synthesized by taking the inverse DCT (I-DCT) of the output $X_{k}$ of the L/NL regression model:
$x_{n}=\sum_{k=0}^{Q-1}\alpha_{k}X_{k}\cos\left(\frac{\pi}{Q}\left(n+\frac{1}{2}\right)k\right)\quad n=0,\ldots,Q-1$ (11)
where the coefficient $\alpha_{k}$ is defined as:
$\alpha_{k}=\begin{cases}\sqrt{\frac{1}{Q}}&\text{if}\ k=0\\ \sqrt{\frac{2}{Q}}&\text{if}\ k>0\end{cases}$ (12)
Fig. 1 provides a pictorial overview of the proposed FD L/NL regression model.